# Scaling Off-Policy Reinforcement Learning with Batch and Weight Normalization

### v1

@@ Abstract
Reinforcement Learning, Batch Normalization, Weight Normalization
@@ 1 Introduction
Different approaches have been explored to address the problem of low sample efficiency in rl. Model-based rl, on the one hand, attempts to increase sample efficiency by learning dynamic models that reduce the need for collecting real data, a process often expensive and time-consuming (Sutton, 1990; Janner et al., 2019; Feinberg et al., 2018; Heess et al., 2015). Model-free rl approaches, on the other hand, have explored increasing the number of gradient updates on the available data, referred to as the update-to-data (utd) ratio (Nikishin et al., 2022; D’Oro et al., 2022), modifying network architectures (Bhatt et al., 2024), or both (Chen et al., 2021; Hiraoka et al., 2021; Hussing et al., 2024; Nauman et al., 2024).
Figure 1: CrossQ + wn utd$\mathbin{=}5$ is competitive to bro utd$\mathbin{=}10$. In comparison, our proposed CrossQ + wn is a simple algorithm that does not require extra exploration policies or full parameter resets. We present results for $25$ complex continuous control tasks from the dmc and MyoSuite benchmarking suites. $1.0$ marks the maximum score achievable on the respective benchmarks (dmc return up to $1000$ / Myosuite up to $100\%$ success rate).
In this work, we build upon CrossQ (Bhatt et al., 2024), a recent model-free rl algorithm that recently showed state-of-the-art sample efficiency on the MuJoCo (Todorov et al., 2012) continuous control benchmarking tasks. Notably, the authors achieved this by carefully utilizing Batch Normalization (bnIoffe (2015)) within the actor-critic architecture. A technique previously thought not to work in rl, as famously reported by Hiraoka et al. (2021) and others. The insight that Bhatt et al. (2024) offered is that one needs to carefully consider the different state-action distributions within the Bellman equation and handle them correctly to succeed. This novelty allowed CrossQ at a low utd of $1$ to outperform the then state-of-the-art algorithms that scaled their utd ratios up to $20$. Even though higher utd ratios are more computationally expensive, they allow for larger policy improvements using the same amount of data.
This naturally raises the question: How can we extend the sample efficiency benefits of CrossQ and bn to the high utd training regime? Which we address in this manuscript.
@@ 2 Preliminaries
A Markov Decision Process (mdp(Puterman, 2014) is a tuple $\mathcal{M}=\langle\mathcal{S},\mathcal{A},\mathcal{P},r,\mu_{0},\gamma\rangle$, with state space $\mathcal{S}\subseteq\mathbb{R}^{n}$, action space $\mathcal{A}\subseteq\mathbb{R}^{m}$, transition probability $\mathcal{P}:\mathcal{S}\times\mathcal{A}\rightarrow{}\Delta(\mathcal{S})$, the reward function $r:\mathcal{S}\times\mathcal{A}\rightarrow{}\mathbb{R}$, initial state distribution $\mu_{0}$ and discount factor $\gamma$. We define the rl problem according to Sutton & Barto (2018). A policy $\pi:\mathcal{S}\rightarrow{}\Delta(\mathcal{A})$ is a behavior plan, which maps a state $\bm{s}$ to a distribution over actions $\bm{a}$. The discounted cumulative return is defined as
$\mathcal{R}(\bm{s},\bm{a})=\sum_{t=0}^{\infty}\gamma^{t}r(\bm{s}_{t},\bm{a}_{t}),$ (1)
where $\bm{s}_{0}=\bm{s}$ and $\bm{a}_{0}=\bm{a}$. Further, $\bm{s}_{t+1}\sim\mathcal{P}(\>\cdot\>|\bm{s}_{t},\bm{a}_{t})$ and $\bm{a}_{t}\sim\pi(\>\cdot\>|\bm{s}_{t})$. The Q-function of a policy $\pi$ is the expected discounted return $Q^{\pi}(\bm{s},\bm{a})=\mathbb{E}_{\pi,\mathcal{P}}[\mathcal{R}(\bm{s},\bm{a})].$
The goal of an rl agent is to find an optimal policy $\pi^{*}$ that maximizes the expected return from the initial state distribution
$\pi^{*}=\textstyle\operatorname*{arg\,max}_{\pi}\mathbb{E}_{\bm{s}\sim\mu_{0}}\left[Q^{\pi}(\bm{s},\bm{a})\right].$ (2)
@@
Soft Actor-Critic (sacHaarnoja et al. (2018)) addresses this optimization problem by jointly learning neural network representations for the Q-function and the policy. The policy network is optimized to maximize the Q-values, while the Q-function is optimized to minimize the Bellmann error
$\mathcal{L}=\mathop{\mathbb{E}}_{\mathcal{D}}\left[\frac{1}{2}\left(Q_{\theta}(\bm{s}_{t},\bm{a}_{t})-\left(r(\bm{s}_{t},\bm{a}_{t})+\gamma\mathop{\mathbb{E}}_{\mathcal{P}}\left[V(\bm{s}_{t+1})\right]\right)\right)^{2}\right],$ (3)
where the value function is computed by taking an expectation over the learned Q function
$V(\bm{s}_{t+1})=\textstyle\mathop{\mathbb{E}}_{\mathcal{P},\pi_{\theta}}\left[Q_{\bar{\theta}}(\bm{s}_{t+1},\bm{a}_{t+1})\right].$ (4)
@@
To stabilize the Q-function learning, Haarnoja et al. (2018) found it necessary to use a target Q-network in the computation of the value function instead of the regular Q-network. The target Q-network is structurally equal to the regular Q-network, and its parameters $\bar{\theta}$ are obtained via Polyak Averaging over the learned parameters $\theta$. While this scheme ensures stability during training by explicitly delaying value function updates, it also arguably slows down online learning (Plappert et al., 2018; Kim et al., 2019; Morales, 2020).
@@
Instead of relying on target networks, CrossQ (Bhatt et al., 2024) addresses training stability issues by introducing Batch Normalization (bnIoffe (2015)) in its Q-function and achieves substantial improvements in sample and computational efficiency over SAC. A central challenge when using bn in Q networks is distribution mismatch: during training, the Q-function is optimized with samples $\bm{s}_{t},\bm{a}_{t}$ from the replay buffer. However, when the Q-function is evaluated to compute the target values (Equation 4), it receives actions sampled from the current policy $\bm{a}_{t+1}\sim\pi_{\theta}(\;\cdot\;|\bm{s}_{t+1})$. Those samples have no guarantee of lying within the training distribution of the Q-function. bn is known to struggle with out-of-distribution samples, as such, training can become unstable if the distribution mismatch is not correctly accounted for (Bhatt et al., 2024). To deal with this issue, CrossQ removes the separate target Q-function and evaluates both Q values during the critic update in a single forward pass, which causes the bn layers to compute shared statistics over the samples from the replay buffer and the current policy. This scheme effectively tackles distribution mismatch problems, ensuring that all inputs and intermediate activations are effectively forced to lie within the training distribution.
@@
Normalization techniques are widely recognized for improving the training of neural networks, as they generally accelerate training and improve generalization (Huang et al., 2023). There are many ways of introducing different types of normalizations into the rl framework. Most commonly, authors have used Layer Normalization (ln) within the network architectures to stabilize training (Hiraoka et al., 2021; Lyle et al., 2024). Recently, CrossQ has been the first algorithm to successfully use bn layers in rl (Bhatt et al., 2024). The addition of bn leads to substantial gains in sample efficiency. In contrast to ln, however, one needs to carefully consider the different state-action distributions within the critic loss when integrating bn. In a different line of work, Hussing et al. (2024) proposed the integration of unit ball normalization and projected the output features of the penultimate layer onto the unit ball in order to reduce Q-function overestimation.
@@
Although scaling up the utd ratio is an intuitive approach to increase the sample efficiency, in practice, it comes with several challenges. Nikishin et al. (2022) demonstrated that overfitting on early training data can inhibit the agent from learning anything later in the training. The authors dub this phenomenon the primacy bias. To address the primacy bias, they suggest to periodically reset the network parameters while retraining the replay buffer. Many works that followed have adapted this intervention (D’Oro et al., 2022; Nauman et al., 2024). While often effective, regularly resetting is a very drastic intervention and by design induces regular drops in performance. Since the agent has to start learning from scratch repeatedly, it is also not very computing efficient. Finally, the exact reasons why parameter resets work well in practice are not yet well understood (Li et al., 2023). Instead of resetting there have also been other types of regularization that allowed practitioners to train stably with high utd ratios. Janner et al. (2019) generate additional modeled data, by virtually increasing the utd. In redq, Chen et al. (2021) leverage ensembles of Q-functions, while Hiraoka et al. (2021) use dropout and ln to effectively scale to higher utd ratios.
Figure 2: Q-bias and weight norms. CrossQ critic weight norms increase significantly with increasing utd ratios.
3 CrossQ fails to scale up stably
@@
Bhatt et al. (2024) demonstrated CrossQ’s state-of-the-art sample efficiency on the MuJoCo task suite (Todorov et al., 2012), while at the same time also being very computationally efficient. However, on the more extensive dmc and Myosuite task suites, we find that CrossQ requires tuning. We further find that it works on some, but not all, environments stably and reliably.
Mixed performance of CrossQ.
Figure 2 shows CrossQ training performance on a subset of dmc tasks. Namely, the dog-stand, dog-trot and humanoid-walk, selected for their varying difficulty levels to demonstrate a wide range of behaviors, including both successful learning and challenges encountered during training. The figure compares a sac baseline with standard hyperparameters against tuned CrossQ agents with utd ratios of $1$ and $5$, where the hyperparameters were identified through a grid search over learning rates and network sizes, as detailed in Table 1. The first row of the figure shows the iqm training performance and $95\%$ confidence intervals for each agent across $10$ seeds. Here, we identify three different training behaviors. On dog-stand CrossQ trains stably at utd$=1$, but increasing the utd to $5$ introduces instabilities and decreases performance. On dog-trot, both utd ratios perform very similarly. Finally, on humanoid-walk, utd$\mathbin{=}5$ outperforms utd$\mathbin{=}1$, although it merely manages to catch up to the sac baseline in this case. Overall, for all CrossQ runs we notice very large confidence intervals.
Q-function bias analysis.
The second row in Figure 2 shows the standard deviation of the normalized Q-function bias. This bias measures how much the Q-function is overestimating or underestimating the true expected return of the current policy.
To compute the normalized Q-function bias, we follow the protocol of Chen et al. (2021). We gather $5$ trajectories in the environment using the current policy and use each trajectory’s first $350$ state-action pairs to calculate the bias. For this, we compare the cumulative discounted rewards for each state-action pair with its Q-value predicted by the Q-function. This bias is then normalized using the cumulative discounted rewards. The mean over these normalized Q-function biases measures the expected bias. Even if the mean is high, as long as the bias is consistent, the selected actions of the policy do not change and, therefore, it is not a problem (Van Hasselt et al., 2016). If the standard deviation is high, the change in bias is high, which might hinder learning. Thus, following the work of Chen et al. (2021); Bhatt et al. (2024), we focus on the standard deviation.
We find large fluctuations for the Q-function bias standard deviation with all three agents across environments. And even on dog-stand, where CrossQ utd$\mathbin{=}5$ does not learn reliably, it maintains a small Q-function bias standard deviation. From these mixed results, we conclude that we cannot directly link the standard deviation of the Q-function bias to the learning performance. While this does not mean that for larger utd ratios, the Q-bias would not explode. Rather, it means that, in our case, the Q-bias is not at fault, and the root cause for unstable training lies elsewhere.
@@
The third row of Figure 2 displays the sum over the L2 norms of the dense layers in the critic network. This includes three dense layers, each with a hidden dimension of $512$.
All three baselines exhibit growing network weights over the course of training. We find that the effect is particularly pronounced for CrossQ with increasing utd ratios. We further find that in the second training phase, the spread of network weight norms increases for the dog runs with large confidence intervals. This is visualized by the large spread of the shaded areas, which show the $95\%$ inter percentile ranges for the weight norms.
Growing network weights have been linked to a loss of plasticity, a phenomenon where networks become increasingly resistant to parameter update, which can lead to premature convergence (Elsayed et al., 2024). Additionally, the growing magnitudes pose a challenge for optimization, connected to the issue of growing activations, which has recently been analyzed by Hussing et al. (2024). Further, growing network weights decrease the effective learning rate when the networks contain normalization layers (Van Hasselt et al., 2019; Lyle et al., 2024).
In summary, the scaling results for vanilla CrossQ are mixed. While increasing utd ratios is known to yield increased sample efficiency, if careful regularization is used (Janner et al., 2019; Chen et al., 2021; Nikishin et al., 2022), CrossQ alone with bn cannot benefit from it. We notice that with increasing utd ratios, CrossQ’s weight layer norms grow significantly faster and overall larger. This observation motivates us to further study the weight norms in CrossQ to increase utd ratios.
4 Combining batch normalization and weight normalization for scaling up
Inspired by the combined insights of Van Hasselt et al. (2019) and Lyle et al. (2024), we propose to integrate CrossQ with Weight Normalization (wn) as a means of counteracting the rapid growth of weight norms we observe with increasing update-to-data (utd) ratios.
Our approach is based on the following reasoning: Due to the use of bn in CrossQ, the critic network exhibits scale invariance, as previously noted by Van Laarhoven (2017).
Theorem 4.1 (Van Laarhoven (2017)).

### v2

@@ 1 Introduction
Different approaches have been explored to address the problem of low sample efficiency in rl. Model-based rl, on the one hand, attempts to increase sample efficiency by learning dynamic models that reduce the need for collecting real data, a process often expensive and time-consuming [38, 21, 11, 16]. Model-free rl approaches, on the other hand, have explored increasing the number of gradient updates on the available data, referred to as the update-to-data (utd) ratio [32, 9], modifying network architectures [4], or both [8, 17, 19, 31]. A central tension in these research directions is balancing the sample efficiency of the agent against the computational complexity, i.e., wall-clock time, of the underlying algorithm. Algorithmic adjustments such as model-based rollouts, computing auxiliary exploration rewards and higher utds can all significantly increase the wall-clock time of the algorithm. Likewise, architectural changes such as larger models [31] and more ensemble members [8] also increase the wallclock time of the method. Ideally, we desire architectural and algorithmic changes that balance sample efficiency and simplicity, such as CrossQ [4], which showed that careful use of batch normalization unlocks significantly greater sampler efficiency of deep actor-critic methods without significantly impacting the wallclock time.
Figure 1: CrossQ + wn utd$\mathbin{=}2$ outperforms simba utd$\mathbin{=}2$ and bro utd$\mathbin{=}2$. In comparison, our proposed CrossQ + wn is a simple algorithm that, unlike bro, does not require extra exploration policies or full parameter resets. We present results for $25$ complex continuous control tasks from the dmc and MyoSuite benchmarking suites. $1.0$ marks the maximum score achievable on the respective benchmarks (dmc return up to $1000$ / MyoSuite up to $100\%$ success rate). We present iqm and $90\%$ stratified bootstrap confidence intervals aggregated over multiple environments and $10$ seeds each.
In this work, we build upon CrossQ [4], the model-free rl algorithm that showed state-of-the-art sample efficiency on the MuJoCo [41] continuous control benchmarking tasks, and also enabled learning omni-directional locomotion policies in 8 minutes of real-world experience [6]. Notably, the authors achieved this by carefully utilizing batch normalization (bnIoffe [20]) within the actor-critic architecture. A technique previously thought not to work in rl, as reported by Hiraoka et al. [17] and others. The insight that Bhatt et al. [4] offered is that one needs to carefully consider the different state-action distributions within the Bellman equation and handle them correctly to succeed. This novelty allowed CrossQ at a low utd of $1$ to outperform the then state-of-the-art algorithms that scaled their utd ratios up to $20$. Even though higher utd ratios are more computationally expensive, they allow for larger policy improvements using the same amount of data. This naturally raises the question: How can we extend the sample efficiency benefits of CrossQ and bn to the high utd training regime? Which we address in this manuscript.
Figure 2: Comparing performance against wall clock time, measured in environment steps per second (so larger is better) on a single RTX 4090 workstation, we observe the CrossQ + WN outperforms simbaand bro across all environments. We present results for $25$ complex continuous control tasks from the dmc and MyoSuite benchmarking suites. $1.0$ marks the maximum score achievable on the respective benchmarks (dmc return up to $1000$ / MyoSuite up to $100\%$ success rate).
@@ 2 Preliminaries
A Markov decision process (mdp[34] is a tuple $\mathcal{M}=\langle\mathcal{S},\mathcal{A},\mathcal{P},r,\mu_{0},\gamma\rangle$, with state space $\mathcal{S}\subseteq\mathbb{R}^{n}$, action space $\mathcal{A}\subseteq\mathbb{R}^{m}$, transition probability $\mathcal{P}:\mathcal{S}\times\mathcal{A}\rightarrow{}\Delta(\mathcal{S})$, the reward function $r:\mathcal{S}\times\mathcal{A}\rightarrow{}\mathbb{R}$, initial state distribution $\mu_{0}$ and discount factor $\gamma$. We define the rl problem according to Sutton and Barto [39]. A policy $\pi:\mathcal{S}\rightarrow{}\Delta(\mathcal{A})$ is a behavior plan, which maps a state $\bm{s}$ to a distribution over actions $\bm{a}$. The discounted cumulative return is defined as $\mathcal{R}^{\pi}(\bm{s},\bm{a})=\sum_{t=0}^{\infty}\gamma^{t}r(\bm{s}_{t},\bm{a}_{t})$, where $\bm{s}_{0}=\bm{s}$, $\bm{a}_{0}=a$ and $\bm{s}_{t+1}\sim\mathcal{P}(\>\cdot\>|\bm{s}_{t},\bm{a}_{t})$ and $\bm{a}_{t}\sim\pi(\>\cdot\>|\bm{s}_{t})$ otherwise. The Q-function of a policy $\pi$ is the expected discounted return $Q^{\pi}(\bm{s},\bm{a})=\mathbb{E}[\mathcal{R}^{\pi}(\bm{s},\bm{a})].$ The goal of an rl agent is to find an optimal policy $\pi^{*}$ that maximizes the expected return from the initial state distribution $\pi^{*}=\textstyle\mathop{\mathrm{arg\,max}}_{\pi}\mathbb{E}_{\bm{s}\sim\mu_{0}}\left[Q^{\pi}(\bm{s},\bm{a})\right].$
@@
Soft actor-critic (sacHaarnoja et al. [13]) addresses this optimization problem by jointly learning neural network representations for the Q-function and the policy. The policy network is optimized to maximize the Q-values, while the Q-function is optimized to minimize the squared Bellmann error, where the value target is computed by taking an expectation over the learned Q function
$V(\bm{s}_{t+1})=\textstyle\mathop{\mathbb{E}}_{\bm{a}_{t+1}\sim\pi_{\theta}(\cdot\mid\bm{s}_{t+1})}\left[Q_{\bar{\theta}}(\bm{s}_{t+1},\bm{a}_{t+1})\right].$ (1)
@@
To stabilize the Q-function learning, Haarnoja et al. [13] found it necessary to use a target Q-network in the computation of the value function instead of the regular Q-network. The target Q-network is structurally equal to the regular Q-network, and its parameters $\bar{\theta}$ are obtained via Polyak Averaging over the learned parameters $\theta$. While this scheme ensures stability during training by explicitly delaying value function updates, it also arguably slows down online learning [33, 22, 30].
@@
Instead of relying on target networks, CrossQ [4] addresses training stability issues by introducing batch normalization (bnIoffe [20]) in its Q-function and achieves substantial improvements in sample and computational efficiency over SAC. A central challenge when using bn in Q networks is distribution mismatch: during training, the Q-function is optimized with samples $\bm{s}_{t},\bm{a}_{t}$ from the replay buffer. However, when the Q-function is evaluated to compute the target values (eq. 1), it receives actions sampled from the current policy $\bm{a}_{t+1}\sim\pi_{\theta}(\;\cdot\;|\bm{s}_{t+1})$. Those samples have no guarantee of lying within the training distribution of the Q-function. bn is known to struggle with out-of-distribution samples, as such, training can become unstable if the distribution mismatch is not correctly accounted for [4]. To deal with this issue, CrossQ removes the separate target Q-function and evaluates both Q values during the critic update in a single forward pass, which causes the bn layers to compute shared statistics over the samples from the replay buffer and the current policy. This scheme effectively tackles distribution mismatch problems, ensuring that all inputs and intermediate activations are effectively forced to lie within the training distribution.
@@
Normalization techniques are widely recognized for improving the training of neural networks, as they generally accelerate training and improve generalization [18]. There are many ways of introducing different types of normalizations into the rl framework. Most commonly, authors have used layer normalization (ln) within the network architectures to stabilize training [17, 29, 26]. Recently, CrossQ has been the first algorithm to successfully use bn layers in rl [4]. The addition of bn leads to substantial gains in sample efficiency. In contrast to ln, however, one needs to carefully consider the different state-action distributions within the critic loss when integrating bn. In a different line of work, Hussing et al. [19] proposed the integration of unit ball normalization and projected the output features of the penultimate layer onto the unit ball in order to reduce Q-function overestimation.
@@
Although scaling up the utd ratio is an intuitive approach to increase the sample efficiency, in practice, it comes with several challenges. Nikishin et al. [32] demonstrated that overfitting on early training data can inhibit the agent from learning anything later in the training. The authors dub this phenomenon the primacy bias. To address the primacy bias, they suggest to periodically reset the network parameters while retraining the replay buffer. Many works that followed have adapted this intervention [9, 31]. While often effective, regularly resetting is a very drastic intervention and by design induces regular drops in performance. Since the agent has to start learning from scratch repeatedly, it is also not very computing efficient. Finally, the exact reasons why parameter resets work well in practice are not yet well understood [27]. Instead of resetting there have also been other types of regularization that allowed practitioners to train stably with high utd ratios. Janner et al. [21] generate additional modeled data, by virtually increasing the utd. In redq, Chen et al. [8] leverage ensembles of Q-functions, while Hiraoka et al. [17] use dropout and ln to effectively scale to higher utd ratios.
3 Batch normalization alone fails to scale with respect to task complexity
Figure 3: Growing parameter norms hinder learning. The performance benefits of CrossQ fail to scale to more complex, higher dimensional tasks such as humanoid locomotion and muscular manipulation. Investigating this, we find that the critic parameter norms increase significantly with increasing utd ratios. As a result, the effective learning rate (ELR) drops and the number of dead neurons increases. Regularizing the critic parameters with weight norm (WN), we successfully mitigate the parameter norm growth and therefore maintain a more consistent ELR, leading to better performance on these more complex tasks. Uncertainty quantification depicts the 90% stratified bootstrap confidence intervals.
@@ 3 Batch normalization alone fails to scale with respect to task complexity
Bhatt et al. [4] demonstrated CrossQ’s state-of-the-art sample efficiency on the MuJoCo task suite [41], while at the same time also being very computationally efficient. However, on the more extensive dmc and MyoSuite task suites, we find that CrossQ requires tuning. We further find that it works on some, but not all, environments stably and reliably.
Figure 3 shows CrossQ training performance on the dmc dog and humanoid and the Myo Hard tasks aggregated by environment suite. We plot the iqm and $90\%$ confidence intervals for each metric across $10$ seeds. The figure compares a CrossQ with utds $1$ and $5$, where the hyperparameters were identified through a grid search over learning rates and network sizes, as detailed in Table 1. The first shows the agent’s training performance, while the other columns show network parameter norms, gradient norms, the effective learning rate [43], and the fraction of dead neurons. Here, we identify three different training behaviors. We notice that CrossQ does not benefit from higher utd ratios, but performance remains similar on the provided tasks. Overall, for all CrossQ runs we notice large confidence intervals.
@@
The second column of Figure 3 displays the sum over the L2 norms of the dense layers in the critic network. This includes three dense layers, each with a hidden dimension of $512$. On both task suites, CrossQ exhibits growing network weights over the course of training. We find that the effect is particularly pronounced for CrossQ with increasing utd ratios. Growing network weights have been linked to a loss of plasticity, a phenomenon where networks become increasingly resistant to parameter update, which can lead to premature convergence [10]. Additionally, the growing magnitudes pose a challenge for optimization, connected to the issue of growing activations, which has recently been analyzed by Hussing et al. [19]. Further, growing network weights decrease the effective learning rate when the networks contain normalization layers [42, 29].
In summary, the scaling results for vanilla CrossQ are mixed. While increasing utd ratios is known to yield increased sample efficiency, if careful regularization is used [21, 8, 32], CrossQ alone with bn cannot benefit from it. We notice that with increasing utd ratios, CrossQ’s weight layer norms grow significantly faster and overall larger. This observation motivates us to further study the weight norms in CrossQ to increase utd ratios.
4 Combining batch normalization and weight normalization enables scaling
Inspired by the combined insights of Van Hasselt et al. [42] and Lyle et al. [29], we propose to integrate CrossQ with weight normalization (Salimans and Kingma [35], wn) as a means of counteracting the rapid growth of weight norms we observe with increasing update-to-data (utd) ratios. A weight normalized parameter $\tilde{\bm{w}}$ is constrained to have an L2 norm of $c$, an additional hyperparameter,
$\tilde{\bm{w}}=c\,\bm{w}\,/\,||\bm{w}||_{2}.$ (2)
Our approach is based on the following reasoning: Due to the use of bn in CrossQ, the critic network exhibits scale invariance, as previously noted by Van Laarhoven [43].
Theorem 1 (Van Laarhoven [43])